If Lilia paid P980.00 in interest at 14% for 2 years, how much money did she borrow?

 


1
Expert's answer
2021-11-17T17:18:19-0500

Solution:

"I=980, R=14,t=2"

Let the money she borrowed be "P"

We know, "I=\dfrac{PRT}{100}"

"\Rightarrow 980=\dfrac{P\times 14\times2}{100}
\\\Rightarrow P=\dfrac{98000}{28}
\\\Rightarrow P=3500"

The money she borrowed was P3500.